# Assist Research

### v1

@@ Abstract
Numerous studies have assessed the proficiency of AI systems, particularly large language models (LLMs), in facilitating everyday tasks such as email writing, question answering, and creative content generation. However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers. In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in four fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not. AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis. An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks. We will release the AAAR-1.0 and keep iterating it to new versions. Project Webpage:   https://renzelou.github.io/AAAR-1.0/
@@ 1 Introduction
To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks. As illustrated in Figure 1, AAAR-1.0 decomposes four distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews. To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering. All four tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed. Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
@@
  • With a random guess baseline of 25%, the performance of most LLMs on EqInfer hovers just slightly above chance, with the top models reaching around 60%. This highlights the difficulty of the task, despite its reliance only on local context reasoning.
@@
  • In ReviewCritique, LLMs struggle to effectively identify deficient human reviews, indicating limited usefulness in assisting meta-reviewers in evaluating the quality of individual paper reviews.
@@ 2 Related Work
With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024). For example, Si et al. (2024) conducted a large-scale human investigation on LLM-generated research ideas and found that LLMs can generate novel ideas compared with humans while lacking feasibility. Du et al. (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper. Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024; 2023; Tang et al., 2023). Notably, Lu et al. (2024) proposed AI-Scientist, an autonomous agent framework that can handle a series of challenging research tasks consecutively, including generating research ideas, coming up with the corresponding experiments along with the implementations, and then writing the final research paper — exactly how human conduct a whole research pipeline. However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task. Our work focuses on building a benchmark that has individual research steps with clear input-output expectations, thus making it suitable for comprehensive LLMs evaluation.
@@
Existing “LLM assists research” benchmarks mainly focus on the implementation and execution part of the research pipeline (Lu et al., 2024; Chen et al., 2024b; Li et al., 2024). For instance, Huang et al. (2024) proposed MLAgentBench to test the LLMs’ capacity for writing project code and training the ML models, where the evaluation metric is the test performance of the models trained by LLMs. However, real-world AI research activities are diverse and some of them are hard to assess for quality, such as generating research ideas, which requires intensive manual assessment (Si et al., 2024; Liang et al., 2024), or LLM-based estimation (Lu et al., 2024). Our work centers on tasks that emphasize a comprehensive mastery of the scientific research field and core elements of a researcher’s daily workload, and we try to build curated task-specific metrics for every single task for a more efficient and accurate LLMs appraisal.
@@
Figure 2 provides an overview of the construction process of AAAR-1.0. In the following sections, we elaborate on the data collection details of the aforementioned four tasks, including § 3.1 EquationInference (EqInfer), § 3.2 ExperimentDesign (ExpDesign), § 3.3 PaperWeakness (Weakness), and § 3.4 ReviewCritique.
@@ 3.1 EquationInference
Crafting a correct scientific equation in paper writing or validating an equation in paper reviewing is challenging, as it requires a thorough understanding of an algorithm or the intricate relationships among numerous variables. Directly prompting LLMs to generate equations proves overly demanding. Therefore, this work formulates EqInfer (Figure 1) as as a multiple-choice equation inference task based on contextual cues, curated as follows.
@@
For the data source, we adopt the pre-compilation LaTeX code for two reasons: i) existing PDF parsing tools, such as PyMuPDF and PaperMage (Lo et al., 2023), can introduce considerable noise to the parsed equation text; ii) considering most of exiting LLMs are capable with processing LaTeX code, using LaTeX source instead of parsed text can be more accurate and provide LLMs with richer information. Meanwhile, to avoid using any low-quality human-written equations, we only crawl those peer-reviewed papers accepted by top-tier conferences. Accordingly, we first obtain the accepted paper list from ACL Anthology, from year 2019 to 2023. Next, we search each paper on arXiv to crawl its LaTeX source (if it exists). Finally, we get a total of 1,762 papers’ source LaTeX packages.
We then clean the LaTeX sources by deleting all the comments and combining multiple cross-referred .tex files into a main file. Afterwards, we use regex to randomly extract (at most) 3 equations’ code snippets per paper, finally resulting in 3,877 human-written equations are extracted.
@@
As we formulate this task as classification, for each human-written positive equation, we have to craft at least three counterpart negative equations. To this end, we prompt GPT-4 to synthesize more equations based on the paper context. For each positive equation, we repeat this prompting (with a high decoding temperature) until three different negative equations are synthesized.
@@
However, the LLM-synthetic equations can sometimes be context-unaligned, i.e., some synthesized equations contain notations that are never defined in the paper context, which is a superficial shortcut for the classification tasks. To improve the data quality, we prompt GPT-4 to identify those context-unaligned negative equations. We then discard those instances where all three negative equations are identified as contextually unaligned. This filtering leads to a final of 1,449 classification instances (62.3% instances are filtered).
@@
Furthermore, it’s also possible that synthesized negative equations are actually correct (i.e., false negative options) — even if the negative and positive equations are written differently, the final compiled results might be the same. To filter out the false negative equations and to have a final check on the classification instances, we then employ human experts to conduct a further data review.
We asked 5 senior Ph.D. students who are experienced in AI research to manually check all the instances. For each classification instance, we ask those human experts to consider the following criteria: i) Are all four equations (both positive and negative) grammatically correct? ii) After compilation, is there only one correct answer? We ask every human expert to use external LaTeX compilation tools (e.g., TeXlive), and identify the instances that cannot meet the criteria. Each instance is examined by at least two experts, and we only keep instances that all experts decide to keep. After this strict examination, a total of 1,049 instances are eventually kept (27.6% instances are filtered)
@@
We finally shuffle the four equations for each classification instance and randomly assign letters (A, B, C, and D) to the equations. We show the data statistics of the final EqInfer in Table 11 and the sample data cases in Figure 7.
@@ 3.2 ExperimentDesign
Intuitively, different experts might have different opinions on the same research topic. Particularly, when explaining the underlying motivation of an experiment, adopting only a single expert’s opinion might introduce bias to our annotation. Hence, we conduct a further multi-round peer discussion. For each online paper PDF, where all the key experiments are identified, summarized, and explained, we ask a different expert (reviewer) to review the annotation by considering the following three criteria: i) Are the identified experiments all the key experiments? ii) Does each experiment summarization covers all key information? iii) Does each explanation sound reasonable and reliable? Each reviewer has to leave comments to the online PDF regarding the above criteria, and then the annotator has to respond to each comment — either accept the suggestion and revise the previous annotation, or provide a “rebuttal” to the reviewer to uphold the annotation. This discussion iterates until both opinions align with each other. Eventually, for each paper, we collect two lists: i) the experiment list, summarizing each experiment step of the paper; ii) the explanation list, the underlying motivations that are one-one corresponding to the experiment.
@@ 4 Evaluation Criteria
For EqInfer, we adopt accuracy as the classification criterion. For ExpDesign and Weakness, since both tasks have natural language outputs, semantic-based metrics are necessary. Hence, in addition to the conventional ROUGE (Lin, 2004), we also develop several novel similarity-based metrics for each specific task.
When evaluating the experiment plan list of ExpDesign, we hope the LLMs can mention as many similar experiment steps as the expert’s plan. Nevertheless, we also don’t expect LLMs to generate too many irrelevant or redundant steps in the plan. This intuition covers both the “recall” and “precision” aspects. Therefore, we develop semantic similarity-based F1 score, denoted as S-F1, which is the harmonic mean of S-Precision and S-Recall:
S-Precision $=\frac{1}{m}\sum_{i=1}^{m}\max_{j}\textrm{sim}(p_{i},g_{j})$ (1)
S-Recall $=\frac{1}{n}\sum_{j=1}^{n}\max_{i}\textrm{sim}(g_{j},p_{i})$ (2)
where the $p$ and $g$ represent the LLM’s prediction plan and the ground-truth plan, respectively. The $m$ and $n$ are the list length of $p$ and $g$ (e.g., $m$ experiment steps in $p$). We use SentenceBERT (Reimers, 2019) to measure the semantic similarity between the $p_{i}$ step and the $g_{j}$ step.
Meanwhile, S-F1 omits the item order difference of two lists, but when giving same-length lists (items have one-one correspondence), we can utilize the following similarity-based matching score:
@@
Unlike ExpDesign, the output of Weakness is multiple reviewers’ weakness lists, which means we have to measure LLM’s single prediction list with a “nested” list. Hence, we rewrite S-Precision ,  S-Recall to SN-Precision , SN-Recall:
SN-Precision $=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{r}\sum_{k=1}^{r}\max_{j}\textrm{sim}(p_{i},g^{k}_{j})\right)$ (4)
SN-Recall $=\frac{1}{r}\sum_{k=1}^{r}\left(\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\max_{i}\textrm{sim}(g^{k}_{j},p_{i})\right)$ (5)
@@
Additionally, in the real world, we would think a review weakness is reliable if it is specific to a paper. Meanwhile, we also hope the review is informative, i.e., no excessive similar weaknesses in one review. Therefore, we adopt the following ITF-IDF metric proposed by Du et al. (2024), which is inspired by the classic TF-IDF:
@@ 5 Experiments and Analyses
In this section, we conduct extensive experiments on AAAR-1.0, across various mainstream LLMs, to quantify the current LLMs’ capacity to tackle high-level research tasks. Specifically, § 5.1 for EqInfer, § 5.2 for ExpDesign, § 5.3 for Weakness, and § 5.4 for ReviewCritique. Please refer to Appendix B.2 for running details of the LLMs.
@@ 5.1 EquationInference
Table 1: Various LLMs’ performances on the 1,049 instances of EqInfer task.
Methods Accuracy (%)
Random Guess 25.00
@@
Gemma 2-27B (Gemma Team, 2024) 3.24
Falcon-40B (Almazrouei et al., 2023) 4.39
OLMo-7B (Groeneveld et al., 2024) 19.00
Mistral-7B (Jiang et al., 2023) 22.21
Qwen 2.5-72B (Qwen Team, 2024) 35.93
Mixtral-8x22B-MoE (Jiang et al., 2024) 37.08
Llama 3.1-70B (MetaAI, 2024) 38.13
@@
Gemini 1.5 Pro (Anil et al., 2023) 34.31
GPT-4o (OpenAI, 2024a) 43.18
GPT-4 (OpenAI et al., 2023) 49.85
o1-preview (OpenAI, 2024b) 59.49
Claude 3.5 sonnet (Anthropic, 2024a) 61.10
Figure 3: The input context length scaling trend on the EqInfer task.
@@
As different LLMs have distinct context windows, to ensure a fair comparison, we fix the maximum input length for all models. According to the data statistics of Table 11, we empirically use 1,000 words for both contexts before and after equations, i.e., 2,000 surrounded words.
@@
Table 1 shows the main results. Firstly, the open-source LLMs, especially the Falcon and Gemma, perform unexpectedly disappointing (even worse than random guesses). These screwed scores are mainly due to the poor long-context instruction following ability, where we find some open-source LLMs are confused with the massive input and often copy the LaTeX code from the input. In contrast, closed-source LLMs generally achieve superior accuracy, probably owing to the richer scientific knowledge from the larger model parameters. However, considering the conventional multi-choice QA formulation of EqInfer, the recently-released GPT-4o solely gets 43.18, implying the unique challenge of EqInfer compared with other scientific QA benchmarks (Song et al., 2023). Notably, with the help of internal CoT, o1 gains stronger performances than GPT-4/GPT-4o, indicating the potential benefits of adopting reasoning for this task.
@@
Table 1 unifies the input context lengths to 1,000 tokens for various LLMs. To answer this question, we experiment with long-context LLMs to investigate the impact of the input context lengths. Particularly, we scale the input length (per side) from 100 to 1,500 words. As shown in Figure 3, for the open-source LLMs (Llama and Qwen), after 300 words length, increasing the input context doesn’t help the performance and even significantly drops Qwen’s scores. While for the closed-source GPT-4-Turbo and GPT-4o, scaling up input length gradually boosts the performances at the first 1,000 words, but stabilizes afterwards. This is in line with human intuition, i.e., surrounding context is required for the equation inference, as the adjacent context usually provides important information, such as the target algorithm description or the notation definition. However, after exceeding a specific threshold, more context information is not beneficial anymore and even confuses those LLMs with poor long-context handling capacity (Wang et al., 2024; Liu et al., 2024).

### v2

# Renze Lou    Hanzi Xu    Sijia Wang    Jiangshu Du    Ryo Kamoi    Xiaoxin Lu    Jian Xie    Yuxuan Sun    Yusen Zhang    Jihyun Janice Ahn    Hongchao Fang    Zhuoyang Zou    Wenchao Ma    Xi Li    Kai Zhang    Congying Xia    Lifu Huang    Wenpeng Yin

@@ Abstract
Numerous studies have assessed the proficiency of AI systems, particularly large language models (LLMs), in facilitating everyday tasks such as email writing, question answering, and creative content generation. However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers. In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in three fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not. AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis. An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks. We will keep iterating AAAR-1.0 to new versions. Project Webpage:   https://renzelou.github.io/AAAR-1.0/
Machine Learning, ICML
[…]
@@ 1 Introduction
To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks. As illustrated in Figure 1, AAAR-1.0 decomposes three distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews. To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering. All three tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed. Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
@@
  • With a random guess baseline of 40% F1, the performance of most LLMs on EqInfer hovers just slightly above chance, with the top models reaching around 46%. This highlights the difficulty of the task, despite its reliance primarily on local context reasoning.
@@
  • In ReviewCritique, LLMs struggle to effectively identify deficient human reviews, indicating limited usefulness in assisting meta-reviewers in evaluating the quality of individual paper reviews. LLMs generated weakness lacks sufficient domain-specific knowledge, model always tends to generate some vague weaknesses that can be applied to any other papers.
@@ 2 Related Work
With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024a; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024; Yu et al., 2024b), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024). For example, Si et al. (2024) conducted a large-scale human investigation on LLM-generated research ideas and found that LLMs can generate novel ideas compared with humans while lacking feasibility. Du et al. (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper. Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024, 2023; Tang et al., 2023). Notably, Lu et al. (2024) proposed AI-Scientist, an autonomous agent framework that can handle a series of challenging research tasks consecutively, including generating research ideas, coming up with the corresponding experiments along with the implementations, and then writing the final research paper — exactly how human conduct a whole research pipeline. However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task. Our work focuses on building a benchmark that has individual research steps with clear input-output expectations, thus making it suitable for comprehensive LLMs evaluation.
@@
Existing “LLM assists research” benchmarks mainly focus on the implementation and execution part of the research pipeline (Lu et al., 2024; Chen et al., 2024a; Li et al., 2024; Chan et al., 2024). For instance, Huang et al. (2024) proposed MLAgentBench to test the LLMs’ capacity for writing project code and training the ML models, where the evaluation metric is the test performance of the models trained by LLMs. However, real-world AI research activities are diverse and some of them are hard to assess for quality, such as generating research ideas, which requires intensive manual assessment (Si et al., 2024; Liang et al., 2024). Our work centers on tasks that emphasize a comprehensive mastery of the scientific research field and core elements of a researcher’s daily workload, and we try to build curated task-specific metrics for every single task for a more efficient and accurate LLMs appraisal.
@@
Figure 2 provides a data construction overview. In the following sections, we elaborate on the data collection details, including § 3.1 EquationInference (EqInfer), § 3.2 ExperimentDesign (ExpDesign), and § 3.3 PaperWeakness (Weakness), and § 3.4 ReviewCritique.
@@ 3.1 EquationInference
Crafting a correct scientific equation in paper writing or validating an equation in paper reviewing is challenging, as it requires a thorough understanding of an algorithm or the intricate relationships among numerous variables. Directly prompting LLMs to generate equations proves overly demanding. Therefore, this work formulates EqInfer (Figure 1) as a binary inference task. [^1] EqInfer also facilitates a multiple-choice QA setting; while we find a binary inference is more challenging for LLMs.
@@
For the data source, we adopt the pre-compilation LaTeX code for two reasons: i) existing PDF parsing tools, such as PyMuPDF and PaperMage (Lo et al., 2023), can introduce considerable noise to the parsed equation text; ii) considering most of exiting LLMs are capable with processing LaTeX code, using LaTeX source instead of parsed text can be more accurate and provide LLMs with richer information. Meanwhile, we only crawl those peer-reviewed papers accepted by top-tier conferences to avoid using low-quality human-written equations. Accordingly, we first obtain the accepted paper list from ACL Anthology, from year 2019 to 2023. Next, we search each paper on arXiv to crawl its LaTeX source (if it exists). Finally, we get a total of 1,762 papers’ source LaTeX packages. We then clean the LaTeX sources by deleting all the comments and combining multiple cross-referred .tex files into a main file. Afterward, we use regex to randomly extract (at most) 3 equations’ code snippets per paper, resulting in 3,877 human-written equations.
@@
As EqInfer assessing whether the LLMs can infer the correctness of equation (i.e., binary classification), for each human-written positive equation, we have to craft counterpart negative equations. To this end, for each positive equation, we prompt GPT-4 to synthesize a negative equation based on the paper context. We repeat this prompt (with a high decoding temperature) until three different negative equations are synthesized. [^2] The number of negative equations is empirically decided.
@@
However, the LLM-synthetic equations can be context-unaligned, i.e., some synthesized equations contain notation that is never defined in the paper context, which becomes a superficial shortcut and too effortless for LLMs to identify. To improve data quality, we prompt GPT-4 to identify context-unaligned negative equations. We then eliminate the positive equation and its negative counterparts, where all three negative counterparts are unaligned. This filtering leads to a final of 1,449 positive equations and 4,347 negative equations (each positive equation has three negative counterparts, and at least one negative counterpart is “challenging”).
@@
Furthermore, it’s also possible that synthesized negative equations are actually correct (i.e., false negative) — even if the negative and positive equations are written differently, the final compiled results might be the same. We then employ human experts to review the data further and filter out false negative equations, checking the classification instances for accuracy.
We asked 5 senior PhD students who are experienced in AI research to check all instances. We ask human experts to consider the following criteria for each positive equation and its negative counterparts (each pair): i) Are all equations grammatically correct? ii) After compilation, are all negative equations different from the positive ones? We ask every human expert to use external LaTeX compilation tools (e.g., TeXlive), and identify the pairs that cannot meet the criteria. Each pair is examined by at least two experts, and we only keep pairs that all experts decide to keep. After this strict examination, a total of 1,049 pairs are eventually kept (27.6% pairs are filtered)
@@
We finally obtain 1,049 positive equations (each has three negative counterparts). We show data statistics of  EqInfer in Table 9 and data examples in Figure 8.
@@ 3.2 ExperimentDesign
Intuitively, different experts might have different opinions on the same research topic. Particularly, when explaining the underlying motivation of an experiment, adopting only a single expert’s opinion might introduce bias to our annotation. Hence, we conduct a further multi-round peer discussion. For each paper, where all the key experiments are identified, summarized, and explained, we ask a different expert (reviewer) to review the annotation by considering the following three criteria: i) Are the identified experiments all the key experiments? ii) Does each experiment summarization covers all key information? iii) Does each explanation sound reasonable and reliable? Each reviewer must leave comments on the online PDF regarding the above criteria, and then the annotator must respond to each comment — either accept the suggestion and revise the previous annotation or provide a “rebuttal” to the reviewer to uphold the annotation. This discussion is iterative until both opinions align. Eventually, for each paper, we collect two lists: i) the experiment list, summarizing each experiment step of the paper; ii) the explanation list, the underlying motivations that are one-one corresponding to the experiment.
@@ 4 Evaluation Criteria
For EqInfer, we adopt F1 as the classification criterion. For ExpDesign and Weakness, since both tasks have free-form outputs, we develop several novel task-specific metrics in addition to the conventional ROUGE (Lin, 2004).
We use LLMs to evaluate the experiment list of ExpDesign. Specifically, given a model-predicted experiment list $p$, and the ground-truth list $g$, we calculate:
En-Precision $=\frac{1}{m}\sum_{i=1}^{m}f(p_{i},g)$ (1)
En-Recall $=\frac{1}{n}\sum_{j=1}^{n}f(g_{j},p)$ (2)
where the $m$ and $n$ are the list length of $p$ and $g$; $f(.)$ represents the LLM prompting, where we prompt LLM to decide whether each predicted experiment item ($p_{i}$) is entailed by the whole ground-truth list ($g$), proceeding with binary output, and vice versa. Intuitively, En-Precision reflects how many prediction experiments match ground-truth experiments. In this work, we used GPT-4o as an evaluator.
While for the explanation generation of ExpDesign, as the prediction experiments are one-on-one corresponding to the ground truth, we adopt a semantic-based metric:
@@
where we use SentenceBERT (Reimers, 2019) to measure the semantic similarity between $p_{i}$ and $g_{j}$.
Unlike ExpDesign, the ground truth of Weakness is multiple reviewers’ weakness lists. Instead of merely merging the opinions of various reviewers into one flattened list and keeping LLM-as-judge as the metric (which is not only costly but also loses the structural information of diverse research perspectives), we employ the following semantic-based metric to efficiently evaluate predicted weaknesses:
S-Precision $=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{r}\sum_{k=1}^{r}\max_{j}\textrm{sim}(p_{i},g^{k}_{j})\right)$ (4)
S-Recall $=\frac{1}{r}\sum_{k=1}^{r}\left(\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\max_{i}\textrm{sim}(g^{k}_{j},p_{i})\right)$ (5)
@@
Additionally, in the real world, we would think a review weakness is reliable if it is specific to a paper. Meanwhile, we also hope the review is informative, i.e., no excessive similar weaknesses in one review. Inspired by the classic TF-IDF, we propose a novel review diversity metric:
@@ 5 Experiments and Analyses
In this section, we conduct extensive experiments on AAAR-1.0, across various mainstream LLMs, to quantify the current LLMs’ capacity to tackle high-level research tasks. Specifically, § 5.1 for EqInfer, § 5.2 for ExpDesign, § 5.3 for Weakness; and § 5.4 for ReviewCritique. Please refer to Appendix B.2 for running details of the LLMs.
@@ 5.1 EquationInference
Table 1: Various LLMs’ performances on EqInfer task (1,049 positive and 3,147 negative samples). “All-positive” indicates a baseline that predicts all equations as positive.
Methods F1 Prec. Rec.
All-Positive 40.00 25.00 100.00
@@
OLMo-7B (Groeneveld et al., 2024) 13.64 11.93 15.91
Mistral-7B (Jiang et al., 2023) 28.45 19.28 54.24
Mixtral-8x22B-MoE (Jiang et al., 2024) 40.90 26.15 93.80
Qwen 2.5-72B (Qwen Team, 2024) 31.22 26.28 57.40
Llama 3.1-70B (MetaAI, 2024) 33.08 22.14 65.39
@@
Gemini 1.5 Pro (Anil et al., 2023) 46.74 32.05 86.27
Claude 3.5 sonnet (Anthropic, 2024a) 45.13 29.48 96.18
GPT-4o (OpenAI, 2024a) 40.35 30.79 58.53
o1-preview (OpenAI, 2024b) 46.35 31.43 88.27
@@
As different LLMs have distinct context windows, to ensure a fair comparison, we fix the maximum input length for all models. According to Table 9, we empirically use 1,000 words for both contexts before and after equations, i.e., 2,000 surrounding words.
@@
Table 1 shows the main results. Firstly, a simple baseline that predicts all equations as positive achieves 40% F1 (due to the 1:3 of positive and negative equations), while nearly all open-source LLMs even cannot beat this naive baseline. Notably, though the performance of Mixtral is slightly superior to the baseline, the extremely biased precision and recall scores imply that Mixtral is also simply predicting almost all samples as positive instead of truly inferring. Meanwhile, compared to the All-Positive baseline, the performance superiority of the strong close-source LLMs is not significant, the best LLM on this task only obtains 46.74%, which demonstrates the challenge of EqInfer compared with other similar benchmarks (Song et al., 2023). The generally high recall with low precision of all LLMs also indicates real-world risks, e.g., relying on LLMs to check the validity of equations in paper review.
@@
EqInfer places high demands on reasoning within the scientific context. To quantify the impact of input context length, we scale the input length (per side) from 100 to 1,500 words. As shown in Figure 4, for the open-source LLMs (Llama and Qwen), an appropriate context length can boost the performance; while for GPT-4o, scaling up the context length doesn’t contribute much to the F1. However, during the scaling, we find that the precision of GPT-4o is gradually increased, and the recall is decreased accordingly; considering the label distribution of EqInfer, we believe precision can better reflect the model’s true capacities on this task. Thus, we anticipate that scaling up context shall be beneficial to those strong close-source LLMs such as GPT-4o.